OpenStudy (anonymous):

MEDAL AND FAN!!!!!! The perimeter of a rectangle is 50 cm and the length is 4 cm longer than the width. Write and solve the system of equations using the substitution method. A. width = 10 cm, length= 14 cm B. width = 10 1/2 cm, length= 14 1/2 cm C. width = 11 cm, length= 15 cm D. width = 23 cm, length= 27 cm

OpenStudy (nincompoop):

ermmmm

OpenStudy (anonymous):

:0

OpenStudy (anonymous):

ummmm okay

OpenStudy (nincompoop):

2L + 2W = P is the formula for perimeter of a rectangle, yes or no?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

There are four sides, so to find the perimeter we just add them all up perimeter=length+length+width+width

OpenStudy (anonymous):

50=2×(width+4)+2×width

OpenStudy (anonymous):

By substituting length with width + 4 we reduced the equation to a single variabale which we can solve.

OpenStudy (anonymous):

so A?

OpenStudy (anonymous):

Mhmm~

OpenStudy (nincompoop):

I think it is asking for the solution to be specifically substitution by system of equation

OpenStudy (anonymous):

.....ihave no clue what you said:(

OpenStudy (nincompoop):

nvm

OpenStudy (nincompoop):

here's what is being sought for: Write and solve the system of equations using the substitution method.

OpenStudy (anonymous):

yeash iknow i just dont know how to

OpenStudy (nincompoop):

that is the purpose of this drill - for you to learn how to do it

OpenStudy (anonymous):

yep

OpenStudy (nincompoop):

2L + 2W = P ------------- 2L + 2W = 50 L = W+4 your composition should start out like this